OpenStudy (anonymous):

cosx/(1-sinX) answers: a. 1-sinx/cosx b. sinx c. 1+secx/secx d. 1+cscx/cscx

OpenStudy (slaaibak):

here is a clue \[{\cos x \over 1 - \sin x} * {1 + \sin x \over 1 + \sin x}\]

OpenStudy (anonymous):

\[cosx+sinxcosx \div 1+sinx ^{2}\] is that right?

OpenStudy (slaaibak):

The denominator should be 1 - sin^2 x Then you can use an identity

OpenStudy (anonymous):

so the answer is sinx right? (: